OpenStudy (anonymous):

In the system –12c + 3d = 6 and –3c + 6d = –9, which expression could you use for substitution

OpenStudy (anonymous):

-12c + 3d = 6 (divide by 3) -4c + d = 2 d = 4c + 2 4c + 2 can be substituted in for d in equation 2 OR -3c + 6d = -9 (divide by -3) c - 2d = 3 c = 2d + 3 2d + 3 can be substituted in for c in equation 1